# Reinforced In-Context Black-Box Optimization

### v1

@@ 3.3 Practical Implementation
Next, we detail the model architecture, training and inference of RIBBO.
Model Architecture. For the formalization of the model $\mathcal{M}_{\bm{\theta}}$, we adopt the commonly used GPT architecture [47], which comprises a stack of causal attention blocks, each containing an attention component and a feed-forward network. We aggregate each triplet $(\bm{x}_{i},y_{i},R_{i})$ using a two-layer MLP network. The output of $\mathcal{M}_{\bm{\theta}}$ is a diagonal Gaussian distribution of the next query point.
Note that previous works that adopt the sequence model as the surrogate model [42; 39; 40] typically remove the positional encoding because the surrogate model should be invariant to the history order. On the contrary, our implementation preserves the positional encoding, naturally following the behavior of certain algorithms (e.g., BO or EA) and making it easier to learn from algorithms. Additionally, the positional encoding can keep the monotonically decreasing order of RTG tokens. More architecture details can be found in Appendix A.
Model Training. RTG tokens are computed as outlined in Eq. (3) for the offline datasets before training. As the calculation of regret requires the optimum value of task $i$, we use the best-observed value $y_{\rm max}^{i}$ as a proxy for the optimum value. Let $\{\widehat{\mathcal{D}}_{i,j}\}_{i\in[N],j\in[K]}$ denote the RTG augmented datasets with $N$ tasks and $K$ algorithms. We sample a minibatch of consecutive subsequences of length $\tau<T$ uniformly from the augmented datasets and train $\mathcal{M}_{\bm{\theta}}$ to minimize the RIBBO loss in Eq. (4).
@@
1:  Initialize context $\hat{\bm{h}}_{0}=\{(\bm{x}_{0},y_{0},R_{0})\}$, where $\bm{x}_{0}$ and $y_{0}$ are placeholders for padding and $R_{0}=0$;
@@
Model Inference. The model $\mathcal{M}_{\bm{\theta}}$ generates the query points $\bm{x}_{t}$ auto-regressively during the inference process, which involves iteratively selecting a new query point $\bm{x}_{t}$ based on the current augmented history $\hat{\bm{h}}_{t-1}$, evaluating the query as $y_{t}=f(\bm{x}_{t})$, and updating the history $\hat{\bm{h}}_{t}=\hat{\bm{h}}_{t-1}\cup\{(\bm{x}_{t},y_{t},R_{t})\}$. A critical aspect of this process is how to specify the value of RTG (i.e., $R_{t}$) at every iteration $t$. Inspired by DT, a naive approach is to specify a desired performance as the initial RTG $R_{0}$, and then decrease it as
$R_{t}=R_{t-1}-(y^{*}-y_{t}).$ (5)
However, this strategy has the risk of producing out-of-distribution RTGs, since the values can fall below 00 due to an improperly selected $R_{0}$.
Given the fact that RTGs are lower bounded by 00 and a 00 value implies a good BBO algorithm with low regret, we propose to set the immediate RTG as 0, and use a strategy called Hindsight Regret Relabelling (HRR) to update previous RTGs based on the current sampling. The inference procedure with HRR is detailed in Algorithm 1. In line 1, the history $\hat{\bm{h}}_{0}$ is initialized with padding placeholders $\bm{x}_{0},y_{0}$ and RTG $R_{0}=0$. At iteration $t$ (i.e., lines 3–7), the model $\mathcal{M}_{\bm{\theta}}$ is first fed with the augmented history $\hat{\bm{h}}_{t-1}$ to generate the next query point $\bm{x}_{t}$ in line 3, and $\bm{x}_{t}$ is evaluated to obtain $y_{t}=f(\bm{x}_{t})$ in line 4. Then, the immediate RTG $R_{t}$ is set to 00, and we employ HRR to update RTG tokens in the history $\hat{\bm{h}}_{t-1}$: calculating the instantaneous regret $r=y^{*}-y_{t}$ in line 5 and adding $r$ to every RTG token in $\hat{\bm{h}}_{t-1}$ in line 6, i.e.,
@@
Note that this relabelling process guarantees that $\forall 0\leq i<t$, the RTG token $R_{i}=\sum_{t^{\prime}=i+1}^{t}(y^{*}-y_{t^{\prime}})$, which can also be written as $\sum_{t^{\prime}=i+1}^{T}(y^{*}-y_{t^{\prime}})$, consistent with the definition in Eq. (3), because the immediate RTG $R_{t}=\sum_{t^{\prime}=t+1}^{T}(y^{*}-y_{t^{\prime}})$ is set to 0. In line 7, the history $\hat{\bm{h}}_{t}$ is updated by expanding $\hat{\bm{h}}_{t-1}$ with $\{(\bm{x}_{t},y_{t},0)\}$, i.e., the current sampling and its immediate RTG $R_{t}=0$. The above process is repeated until reaching the budget $T$. Thus, we can find that HRR not only exploits the full potential of $\mathcal{M}_{\bm{\theta}}$ through using zero as the immediate RTG and thereby demands the model to generate the most advantageous decisions, but also preserves the calculation of RTG tokens following the same way as the training data, i.e., representing the cumulative regret over the future optimization history.
@@ 3.4 Data Generation
Data Collection. Given a set of tasks $\{f_{i}\}_{i=1}^{N}$ sampled from the task distribution $P(\mathcal{F})$, we can select a diverse set of algorithms for data collection. For example, we can select some representatives of different types of BBO algorithms (e.g., BO and EA). Datasets $\mathcal{D}_{i,j}$ are then obtained by using each behavior algorithm to optimize each function with different random seeds. Each optimization history $\bm{h}_{T}=\{(\bm{x}_{t},y_{t})\}_{t=1}^{T}$ in $\mathcal{D}_{i,j}$ is then augmented using RTG tokens $R_{t}$ computed as in Eq. (3). The resulting augmented histories $\hat{\bm{h}}_{T}=\{(\bm{x}_{t},y_{t},R_{t})\}_{t=0}^{T}$ compose the final datasets $\widehat{\mathcal{D}}_{i,j}$ for training the model $\mathcal{M}_{\bm{\theta}}$.
Data Normalization. To provide a unified interface and balance the statistic scales across tasks, it is important to apply normalization to the inputs to our model. We normalize the point $\bm{x}$ by $(\bm{x}-\bm{x}_{\rm min})/(\bm{x}_{\rm max}-\bm{x}_{\rm min})$, with $\bm{x}_{\rm max}$ and $\bm{x}_{\rm min}$ being the upper and lower bounds of the search space, respectively. For the function value $y$, we apply random scaling akin to previous works [61; 14]. That is, when sampling a history $\bm{h}_{\tau}$ from the datasets $\mathcal{D}_{i,j}$, we randomly sample the lower bound $l\sim\mathcal{U}(y^{i}_{\rm min}-\frac{s}{2},y^{i}_{\rm min}+\frac{s}{2})$ and the upper bound $u\sim\mathcal{U}(y^{i}_{\rm max}-\frac{s}{2},y^{i}_{\rm max}+\frac{s}{2})$, where $\mathcal{U}$ stands for uniform distribution, $y^{i}_{\rm min},y^{i}_{\rm max}$ denote the observed minimum and maximum values for $f_{i}$, and $s=y^{i}_{\rm max}-y^{i}_{\rm min}$; the values $y_{t}$ in $\bm{h}_{\tau}$ are then normalized by $(y_{t}-l)/(u-l)$. The RTG tokens are calculated accordingly with the normalized values. The random normalization can make a model exhibit invariance across various scales of $y$.
@@ 4 Experiments
In this section, we examine the performance of RIBBO on a wide range of tasks, including synthetic functions, Hyper-Parameter Optimization (HPO) and robot control problems. The model architecture and hyper-parameters are maintained consistently across these problems. We train our model using five distinct random seeds, ranging from 00 through $4$, and each trained model is also run five times independently during the execution phase. We will report the average performance and standard deviation. Details of the model hyper-parameters can be found in Appendix A. Our code is available at https://github.com/songlei00/RIBBO.
@@ 4.1 Experimental Setup
Benchmarks. We use BBO Benchmarks BBOB [15], HPO-B [2], and rover trajectory planning tasks [57]. The BBOB suite, a comprehensive and widely used benchmark in the continuous domain, consists of $24$ synthetic functions. For each function, a series of transformations are implemented on the search space to obtain a distribution of functions with similar properties. According to the properties of the functions, they can be further divided into $5$ categories, and we select one from each category due to resource constraints, including Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge. HPO-B is the most commonly used HPO benchmark and consists of a series of HPO problems. It fits an XGBoost model as the objective function for each HPO problem, and we conduct experiments on two widely used machine models, SVM and XGBoost. For robot control optimization, we perform experiments on rover trajectory planning, which is a trajectory optimization task to emulate rover navigation. Similar to [15; 56], we implement random translations and scalings to the search space to construct a distribution of functions. For BBOB and rover problems, we sample a set of functions as training and test tasks, while for HPO-B, we use the meta-training/test splits provided by the authors. Detailed explanations of the benchmarks can be found in Appendix B.1.
Data. Similar to OptFormer [14], we employ seven behavior algorithms, i.e., Random Search, Shuffled Grid Search, Hill Climbing, Regularized Evolution [49], Eagle Strategy [63], CMA-ES [26], and GP-EI [8], which are representatives of heuristic search, EA, and BO, respectively. Datasets are generated by using each behavior algorithm to optimize each training function with different random seeds. More details about the behavior algorithms and datasets can be found in Appendix B.2 and B.3.
@@ 4.2 Baselines
As our proposed method RIBBO is an in-context E2E model, the most related baselines are those also training an E2E model with an offline dataset, including Behavior Cloning (BC) [7] and OptFormer [14]. Their hyper-parameters are set as same as that of our model for fairness. Note that the seven behavior algorithms used to generate datasets are also important baselines, which are included for comparison as well.
@@
OptFormer employs transformer to imitate the behaviors of a set of algorithms and needs to manually specify the imitated algorithm during execution. Its original implementation is built upon a text-based transformer with a large training scale. In this paper, we re-implemented a simplified version of OptFormer where we only maintain the algorithm type within the metadata, i.e., employing different initial states $\bm{x}_{0}$ and $y_{0}$ for different algorithms. More details about the training procedure and imitation capacity of the reimplementation can be found in Appendix C.
@@ 4.3 Main Results
RIBBO does not perform well on SVM problem, which may be due to the problem’s low-dimensional nature (only three parameters) and its relative simplicity for optimization. Behavior algorithms can easily achieve good performance, while the complexity of RIBBO’s training and inference processes could instead result in the performance degradation. Regarding to rover problem where GP-EI performs the best, we collect less data from GP-EI than other behavior algorithms due to the high time cost. This may limit RIBBO’s capacity to leverage the high-quality data from GP-EI, given its small proportion relative to the data collected from other behavior algorithms. Despite this, RIBBO is still the runner-up, significantly surpassing the other behavior algorithms.
@@
(b) Cross-dist generalization
@@
Figure 3: (a) Visualization of the contour lines of 2D Branin function and sampling points of RIBBO (red), Eagle Strategy (orange), and Random Search (gray), where the arrows represent the optimization trajectory of RIBBO. (b) Cross-distribution generalization by training the model on $4$ synthetic functions and testing on the remaining one, GriewankRosenbrock, which has different properties. Note that the legend shares with that of Figure 2. (c) Initial RTG $R_{0}$’s influence on performance. (d) RTG update strategy comparison between HRR and the naive strategy with different $R_{0}$.
Cross-Distribution Generalization. We further conduct experiments to examine the generalization of RIBBO to unseen function distributions. For this purpose, we train the model on $4$ of the $5$ chosen synthetic functions and test it with the remaining one. Note that each function (i.e., Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge) here actually represents a distribution of functions with similar properties, and we sample a set of functions from each distribution as introduced before. Greiwank Rosenbrock is used for testing. The results in Figure 3(b) suggest that RIBBO has a strong generalizing ability to unseen function distributions. More results on cross-distribution generalization are deferred to Appendix D.
The good generalization of RIBBO can be attributed to the paradigm of learning the entire algorithm, which can acquire general knowledge, such as exploration and exploitation trade-off from data, as observed in [13]. In contrast, such generalization may be limited if we learn surrogate models from data, because the function landscape inherent to surrogate models will contain only the knowledge of similar functions.
@@ 4.4 Ablation Studies
Influence of Initial RTG Token $R_{0}$. By incorporating RTG tokens into modeling, RIBBO should be able to attend to the RTGs and generate optimization trajectories that achieve the specified initial RTG token $R_{0}$. To validate this, we examine the performance of RIBBO when specifying different values of $R_{0}$, and the results are presented in Figure 3(c). Here, the RTG values, instead of the normalized objective values, of the trajectory are used as the $y$-axis. We can observe that the cumulated regrets of the generated query sequence do correlate with the specified RTG, indicating that RIBBO establishes the connection between regret and generation.
Effectiveness of HRR. A key point of the inference procedure is how to update the value of RTG tokens at each iteration. To assess the effectiveness of the proposed strategy HRR in Eq. (6), we compare it with the naive strategy in Eq. (5), which sets an initial RTG token $R_{0}$ and decreases it by the one-step regret after each iteration. The results are shown in Figure 3(d). The naive strategy displays distinct behaviors when $R_{0}=0$ or $R_{0}=5$. When $R_{0}=0$ i.e. the lower bound of regret, the model performs well in the initial stage. However, as the optimization progresses, the RTG tokens gradually decrease to negative values, leading to poor performance since negative RTGs are out-of-distribution values. Using $R_{0}=5$ compromises the initial performance, as the model may not choose the most aggressive solutions with a high $R_{0}$. However, higher $R_{0}$ yields better convergence value since it prevents out-of-distribution RTGs in the late stage. The proposed HRR strategy performs the best across the whole optimization stage, because setting the immediate RTG as 00 encourages the model to make the most advantageous decisions at every iteration, while hindsight relabeling of previous RTG tokens in Eq. (6) keeps the values meaningful and feasible.
Other Ablation Studies. We also study the effect of the method to aggregate $(\bm{x}_{i},y_{i},R_{i})$ tokens, the normalization method for the function value $y$, the model size, and the length $\tau$ of sampled subsequences. Due to space limitation, we provide the results in Appendix E.

### v2

@@ 3.2 Practical Implementation
Next, we detail the model architecture, training, and inference of RIBBO.
Model Architecture. For the formalization of the model $\mathcal{M}_{\bm{\theta}}$, we adopt the commonly used GPT architecture [48], which comprises a stack of causal attention blocks, each containing an attention component and a feed-forward network. We aggregate each triplet $(\bm{x}_{i},y_{i},R_{i})$ using a two-layer MLP network. The output of $\mathcal{M}_{\bm{\theta}}$ is a diagonal Gaussian distribution of the next query point. Note that previous works that adopt the sequence model as a surrogate model [40, 41, 43] typically remove the positional encoding because the surrogate model should be invariant to the history order. On the contrary, our implementation preserves the positional encoding, naturally following the behavior of certain algorithms (e.g., BO or EA) and making it easier to learn from algorithms. Additionally, the positional encoding can keep the monotonically decreasing order of RTG tokens. More architecture details can be found in Appendix A.
Model Training. RTG tokens are computed as outlined in Eq. (3) for the offline datasets before training. As the calculation of regret requires the optimum value of task $i$, we use the best-observed value $y_{\rm max}^{i}$ as a proxy for the optimum. Let $\{\widehat{\mathcal{D}}_{i,j}\}_{i\in[N],j\in[K]}$ denote the RTG augmented datasets with $N$ tasks and $K$ algorithms. We sample a minibatch of consecutive subsequences of length $\tau<T$ uniformly from the augmented datasets and train $\mathcal{M}_{\bm{\theta}}$ to minimize the RIBBO loss in Eq. (4).
@@
1:  Initialize $\hat{\bm{h}}_{0}=\{(\bm{x}_{0},y_{0},R_{0})\}$, where $\bm{x}_{0}$ and $y_{0}$ are placeholders for padding and $R_{0}=0$;
@@
Model Inference. The model $\mathcal{M}_{\bm{\theta}}$ generates the query points $\bm{x}_{t}$ auto-regressively during the inference process, which involves iteratively selecting a new query point $\bm{x}_{t}$ based on the current augmented history $\hat{\bm{h}}_{t-1}$, evaluating the query as $y_{t}=f(\bm{x}_{t})$, and updating the history $\hat{\bm{h}}_{t}=\hat{\bm{h}}_{t-1}\cup\{(\bm{x}_{t},y_{t},R_{t})\}$. A critical aspect of this process is how to specify the value of RTG (i.e., $R_{t}$) at every iteration $t$. Inspired by DT, a naive approach is to specify a desired performance as the initial RTG $R_{0}$, and decrease it as $R_{t}=R_{t-1}-(y^{*}-y_{t})$. However, this strategy has the risk of producing out-of-distribution RTGs, since the values can fall below 00 due to an improperly selected $R_{0}$.
Given the fact that RTGs are lower bounded by 00 and a 00 value implies a good BBO algorithm with low regret, we propose to set the immediate RTG as 0, and use a strategy called Hindsight Regret Relabelling (HRR) to update previous RTGs based on the current sampling. The inference procedure with HRR is detailed in Algorithm 1. In line 1, the history $\hat{\bm{h}}_{0}$ is initialized with padding placeholders $\bm{x}_{0},y_{0}$ and RTG $R_{0}=0$. At iteration $t$ (i.e., lines 3–7), the model $\mathcal{M}_{\bm{\theta}}$ is fed with the augmented history $\hat{\bm{h}}_{t-1}$ to generate the next query point $\bm{x}_{t}$ in line 3, which is evaluated to obtain $y_{t}$ in line 4. Then, the immediate RTG $R_{t}$ is set to 00, and we employ HRR to update RTG tokens in $\hat{\bm{h}}_{t-1}$, i.e., calculate the instantaneous regret $r=y^{*}-y_{t}$ (line 5) and add $r$ to every RTG token in $\hat{\bm{h}}_{t-1}$ (line 6):
@@
Note that this relabelling process guarantees that $\forall 0\leq i<t$, the RTG token $R_{i}=\sum_{t^{\prime}=i+1}^{t}(y^{*}-y_{t^{\prime}})$, which can also be written as $\sum_{t^{\prime}=i+1}^{T}(y^{*}-y_{t^{\prime}})$, consistent with the definition in Eq. (3), because the immediate RTG $R_{t}=\sum_{t^{\prime}=t+1}^{T}(y^{*}-y_{t^{\prime}})$ is set to 0. In line 7, the history $\hat{\bm{h}}_{t}$ is updated by expanding $\hat{\bm{h}}_{t-1}$ with $\{(\bm{x}_{t},y_{t},0)\}$, i.e., the current sampling and its immediate RTG $R_{t}=0$. This process is repeated until reaching the budget $T$. Thus, we can find that HRR not only exploits the full potential of $\mathcal{M}_{\bm{\theta}}$ through using 00 as the immediate RTG and thereby demands the model to generate the most advantageous decisions, but also preserves the calculation of RTG tokens following the same way as the training data, i.e., representing the cumulative regret over future optimization history.
@@ 3.3 Data Generation
Data Collection. Given a set of tasks $\{f_{i}\}_{i=1}^{N}$ sampled from the task distribution $P(\mathcal{F})$, we can select a diverse set of algorithms for data collection. For example, we can select some representatives from different types of BBO algorithms, e.g., BO and EA. Datasets $\mathcal{D}_{i,j}$ are obtained by using each behavior algorithm to optimize each function with different random seeds. Each optimization history $\bm{h}_{T}=\{(\bm{x}_{t},y_{t})\}_{t=1}^{T}$ in $\mathcal{D}_{i,j}$ is then augmented with RTG tokens $R_{t}$ computed as in Eq. (3). The resulting histories $\hat{\bm{h}}_{T}=\{(\bm{x}_{t},y_{t},R_{t})\}_{t=0}^{T}$ compose the final datasets $\widehat{\mathcal{D}}_{i,j}$ for model training.
Data Normalization. To provide a unified interface and balance the statistic scales across tasks, it is important to apply normalization to the inputs to our model. We normalize the point $\bm{x}$ by $(\bm{x}-\bm{x}_{\rm min})/(\bm{x}_{\rm max}-\bm{x}_{\rm min})$, with $\bm{x}_{\rm max}$ and $\bm{x}_{\rm min}$ being the upper and lower bounds of the search space, respectively. For the function value $y$, we apply random scaling akin to previous works [14, 64]. That is, when sampling a history $\bm{h}_{\tau}$ from the datasets $\mathcal{D}_{i,j}$, we randomly sample the lower bound $l\sim\mathcal{U}(y^{i}_{\rm min}-\frac{s}{2},y^{i}_{\rm min}+\frac{s}{2})$ and the upper bound $u\sim\mathcal{U}(y^{i}_{\rm max}-\frac{s}{2},y^{i}_{\rm max}+\frac{s}{2})$, where $\mathcal{U}$ stands for uniform distribution, $y^{i}_{\rm min},y^{i}_{\rm max}$ denote the observed minimum and maximum values for $f_{i}$, and $s=y^{i}_{\rm max}-y^{i}_{\rm min}$; the values $y_{t}$ in $\bm{h}_{\tau}$ are then normalized by $(y_{t}-l)/(u-l)$. The RTG tokens are calculated accordingly with the normalized values. The random normalization can make a model exhibit invariance across various scales of $y$.
@@ 4 Experiments
In this section, we examine the performance of RIBBO on a wide range of tasks, including synthetic functions, Hyper-Parameter Optimization (HPO) and robot control problems. The model architecture and hyper-parameters are maintained consistently across these problems. We train our model using five distinct random seeds, ranging from 00 to $4$, and each trained model is run five times independently during the execution phase. We will report the average performance and standard deviation. Details of the model hyper-parameters are given in Appendix A. The codes are provided in the supplementary.
@@ 4.1 Experimental Setup
Benchmarks. We use BBO Benchmarks BBOB [15], HPO-B [2], and rover trajectory planning tasks [60]. The BBOB suite, a comprehensive and widely used benchmark in the continuous domain, consists of $24$ synthetic functions. For each function, a series of transformations are implemented on the search space to obtain a distribution of functions with similar properties. According to the properties of the functions, they can be divided into $5$ categories, and we select one from each category due to resource constraints, including Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge. HPO-B is the most commonly used HPO benchmark and consists of a series of HPO problems. It fits an XGBoost model as the objective function for each HPO problem, and we conduct experiments on two widely used machine models, SVM and XGBoost. For robot control optimization, we perform experiments on rover trajectory planning, which is a trajectory optimization task to emulate rover navigation. Similar to [15, 59], we implement random translations and scalings to the search space to construct a distribution of functions. For BBOB and rover problems, we sample a set of functions as training and test tasks, while for HPO-B, we use the meta-training/test splits provided by the authors. Detailed explanations of the benchmarks can be found in Appendix B.1.
Data. Similar to OptFormer [14], we employ $7$ behavior algorithms, i.e., Random Search, Shuffled Grid Search, Hill Climbing, Regularized Evolution [50], Eagle Strategy [66], CMA-ES [26], and GP-EI [8], which are representatives of heuristic search, EA, and BO, respectively. Datasets are generated by using each behavior algorithm to optimize each training function with different random seeds. More details about the behavior algorithms and datasets can be found in Appendix B.2 and B.3.
@@ 4.2 Baselines
As RIBBO is an in-context E2E model, the most related baselines are those also training an E2E model with offline datasets, including Behavior Cloning (BC) [7] and OptFormer [14]. Their hyper-parameters are set as same as that of our model for fairness. Note that the seven behavior algorithms used to generate datasets are also important baselines, and included for comparison as well.
@@
OptFormer employs a transformer to imitate the behaviors of a set of algorithms and needs to manually specify the imitated algorithm during execution. Its original implementation is built upon a text-based transformer with a large training scale. In this paper, we re-implement a simplified version of OptFormer where we only retain the algorithm type within the metadata. Various initial states, denoted as $\bm{x}_{0}$ and $y_{0}$, are used to distinguish between algorithms. They are obtained by indexing the algorithm type through an embedding layer, thereby enabling the identification of distinct behavior algorithms within OptFormer. Further training details can be found in Appendix C.
@@ 4.3 Main Results
RIBBO does not perform well on the SVM problem, which may be due to the problem’s low-dimensional nature (only three parameters) and its relative simplicity for optimization. Behavior algorithms can achieve good performance easily, while the complexity of RIBBO’s training and inference processes could instead result in the performance degradation. For the rover problem where GP-EI performs the best, we collect less data from GP-EI than other behavior algorithms due to the high time cost. This may limit RIBBO’s capacity to leverage the high-quality data from GP-EI, given its small proportion relative to the data collected from other behavior algorithms. Despite this, RIBBO is still the runner-up, significantly surpassing the other behavior algorithms.
Cross-Distribution Generalization. We further conduct experiments to examine the generalization of RIBBO to unseen function distributions. For this purpose, we train the model on $4$ of the $5$ chosen synthetic functions and test it with the remaining one. Note that each function (i.e., Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge) here actually represents a distribution of functions with similar properties, and we sample a set of functions from each distribution as introduced before. Greiwank Rosenbrock is used for testing. The results in Figure 3(b) suggest that RIBBO has a strong generalizing ability to unseen function distributions. More results on cross-distribution generalization are deferred to Appendix D. The good generalization of RIBBO can be attributed to the paradigm of learning the entire algorithm, which can acquire general knowledge, such as exploration and exploitation trade-off from data, as observed in [13]. In contrast, such generalization may be limited if we learn surrogate models from data, because the function landscape inherent to surrogate models will contain only the knowledge of similar functions.
@@
(b) Cross-dist performance
@@
Figure 3: (a) Visualization of the contour lines of $2$D Branin function and sampling points of RIBBO (red), Eagle Strategy (orange), and Random Search (gray), where the arrows represent the optimization trajectory of RIBBO. (b) Cross-distribution generalization by training the model on $4$ synthetic functions and testing on the remaining one, GriewankRosenbrock, which has different properties. The legend shares with that of Figure 2. (c) Initial RTG $R_{0}$’s influence on performance. (d) RTG update strategy comparison between HRR and the naive strategy with various $R_{0}$.
@@ 4.4 Ablation Studies
Influence of Initial RTG Token $R_{0}$. By incorporating RTG tokens, RIBBO can attend to RTGs and generate optimization trajectories based on the specified initial RTG token $R_{0}$. To validate this, we examine the performance of RIBBO with different values of $R_{0}$, and the results are presented in Figure 3(c). Here, the RTG values, instead of normalized objective values, are used as the $y$-axis. We can observe that the cumulative regrets of the generated query sequence do correlate with the specified RTG, indicating that RIBBO establishes the connection between regret and generation.
Effectiveness of HRR. A key point of the inference procedure is how to update the value of RTG tokens at each iteration. To assess the effectiveness of the proposed strategy HRR in Eq. (5), we compare it with the naive strategy, that sets an initial RTG token $R_{0}$ and decreases it by the one-step regret after each iteration. The results are shown in Figure 3(d). The naive strategy displays distinct behaviors when $R_{0}=0$ or $R_{0}=5$. When $R_{0}=0$, i.e., the lower bound of regret, the model performs well initially. However, as the optimization progresses, the RTG tokens gradually decrease to negative values, leading to poor performance since negative RTGs are out-of-distribution values. Using $R_{0}=5$ compromises the initial performance, as the model may not select the most aggressive solutions with a high $R_{0}$. However, a higher $R_{0}$ yields better convergence value since it prevents out-of-distribution RTGs in later stage. The proposed HRR strategy performs the best across the whole optimization stage, because setting the immediate RTG to 00 encourages the model to make the most advantageous decisions at every iteration, while hindsight relabeling of previous RTG tokens in Eq. (5) ensures the values remain meaningful and feasible.
Further Studies. We also study the effects of the method to aggregate $(\bm{x}_{i},y_{i},R_{i})$ tokens, the normalization method for $y$, the model size, and the sampled subsequence length $\tau$. Please see Appendix E. Generalization studies across the entire BBOB suite are shown in Appendix D, while the effects of transformations applied to the search space are detailed in Appendix F.